OpenStudy (anonymous):

x to the 2/7 (power)*x to the 4/5(power)

OpenStudy (anonymous1):

Use the following property: \[x^a*x^b=x^{a+b}\]

OpenStudy (anonymous):

x 8/35?

jhonyy9 (jhonyy9):

x^(2/7) x^(4/5) - so what you need to do with these ?

jhonyy9 (jhonyy9):

there is the first term time second ?

jhonyy9 (jhonyy9):

- so if yes than you need to assum these exponents

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Answer-->x6/12=x1/6

jhonyy9 (jhonyy9):

2/7 +4/5 = ?

jhonyy9 (jhonyy9):

you need to get the denominator common for 7 and 5 - so what will be ?